# references.bib

### v1

@@ 4.6 Visualizations and Clinical Translation
Figure 6: Qualitative results of zero-shot diagnosis on the Rad-ChestCT [draelos2021machine] and BraTS [baid2021rsna, labella2023asnrmiccai] datasets. The first row shows the original clinical scans with pathologic regions outlined (dashed red). The second row shows activation HLIP maps [chefer2021generic]. HLIP identifies the pathologic regions across multiple groups of adjacent chest CT slides (left) and brain MRI scans (right).
@@
Visualizations. In Figure 6, we visualize the activation maps [chefer2021generic] of HLIP in the zero-shot setting on Rad-ChestCT [draelos2021machine] and BraTS [baid2021rsna, labella2023asnrmiccai]. For chest CT (curated, single scan), HLIP attends to pathologic regions across different slices, e.g., slice Z and Z+8, by leveraging scan attention. For brain MRI (uncurated, multiple scans), HLIP attends to pathologic regions across different scan types, e.g., FLAIR and T1W+contrast, by leveraging study attention. This demonstrates the effectiveness HLIP in modeling the hierarchical structure of 3D medical imaging.
@@
Figure 7: Prospective testing results on (a) 52 brain MRI diagnoses and (b) 83 head CT diagnoses. Task-wise results are presented in Figures 12 and 13 in Appendix.
Clinical Translation of HLIP. We conduct a 1-year, health system-scale, comprehensive evaluation on a prospective set of neurological patients. The datasets include ~23K brain MRI studies covering 52 diagnoses and ~15K head CT studies covering 83 diagnoses. We report macro AUC in Figure 7(a) and Figure 7(b). HLIP consistently outperforms ViT. We provide detailed results in Appendix D.
@@ 5 Conclusion
In this work, we introduce HLIP, a scalable language-image pre-training framework for 3D medical imaging. By leveraging the inherent hierarchical structure of radiology studies, HLIP enables pre-training directly on uncurated CT and MRI datasets at health system scale. Across diverse modalities and anatomical regions, including chest CT, head CT, and brain MRI, HLIP achieves state-of-the-art performance on multiple benchmarks. These results demonstrate that, with HLIP, directly pre-training on uncurated clinical datasets is a scalable and effective direction for language-image pre-training in 3D medical imaging. We hope our work facilitates the language-image pre-training in other health systems and inspires future research on scalable approaches for real-world clinical data.
[…]
@@
  • C: Additional Ablation Studies for Chest CT
  • D: Prospective Evaluation
  • E: Complexity
  • F: Discussion
@@ Appendix A Dataset
As mentioned in Section 3.2, this section provides additional details on the two datasets we collected: BrainMRI220K and HeadCT240K. For other datasets, such as CT-RATE [hamamci2024developing], Rad-ChestCT [draelos2021machine], BraTS 2023 [menze2014multimodal, bakas2017advancing, baid2021rsna, labella2023asnrmiccai, moawad2024brain, kazerooni2024brain], NYU-Mets [link2024longitudinal], UCSF-Mets [rudie2024university], RSNA [flanders2020construction], and CQ500 [chilamkurthy2018deep], we refer readers to their original publications. Additionally, we include a discussion of our preprocessing strategy.
@@ A.3 Discussion
To the best of our knowledge, this is the first work to handle 3D medical datasets at the scale of BrainMRI220K and HeadCT240K, which comprise millions of scans. Our preprocessing strategy, introduced in Section 3.2, follows standard practices where possible, but also diverges due to the scale and diversity of our dataset. Specifically, percentile clipping for brain MRI and HU value truncation for head CT follow standard practices [ma2024segment, zhu20253d]. For head CT, we do not concatenate scans with different HU truncations along the channel dimension, as our method is capable of processing multiple scans simultaneously. However, our strategy regarding orientation and spacing may differ from established preprocessing pipelines [isensee2021nnu, blankemeier2024merlin, hamamci2024developing, shui2025large]. Here, we provide a discussion to justify the rationale behind our approach.
@@
Orientation. CT and MRI scans are acquired using different planes, including axial, coronal, and sagittal. Isotropic data are rare in routine radiology studies; the through-plane spacing is typically larger than the in-plane spacing. In the context of batch construction, where all samples in a batch must share the same shape, standardizing the orientation can lead to downsampling along the in-plane axes and upsampling along the through-plane axis. Upsampling in pixel space does not introduce new information, whereas downsampling results in information loss. From the perspective of data augmentation, the orientation can change when 3D random rotation is applied [isensee2021nnu]. So, why do we standardize the orientation at the beginning then augment it during training? Instead, we consistently align the first dimension (depth) with the through-plane axis. Please note that the in-plane orientation is still standardized as RP, RI, or PI. We treat the diversity of acquisition planes (as shown in Figure 8(c) and Figure 10(c)) in our dataset as a form of natural data augmentation.
@@
Spacing. Standardizing spacing across the entire dataset is a common practice in medical image segmentation to ensure that convolution filters interpret anatomical structures consistently [isensee2021nnu]. However, HyperSpace [joutard2024hyperspace] demonstrates that, with appropriate spacing augmentation, convolutional networks can learn spacing variations. Therefore, architectures like Vision Transformers, which can learn scale-equivariant features [dosovitskiy2020image], should also be capable of learning spacing-equivariant features with sufficient spacing augmentation. This is because spacing equivariance is a direct extension of scale equivariance from 2D to 3D space. Given the diversity of spacing settings (as shown in Figure 8(b) and Figure 10(b)) in our dataset, we believe it naturally occupies sufficient spacing augmentation for the model to learn spacing-equivariant features. Therefore, when constructing a batch, we first resize each volume to (48,256,256), and then apply a center crop to obtain (48,224,224).
@@
$M$ scans per study. As the number of scans per study varies, we randomly sample $M=10$ scans per study at each training step to construct a batch. An equal number of positional embeddings is also sampled for these $M$ scans from a total of $M_{\text{max}}=40$ positional embeddings at each step. This means the model is exposed to $M_{\text{max}}$ different positions over the course of training, while observing $M$ positions at each step. This is analogous to a dropout strategy, with a scan dropout rate of $1-\frac{M}{M_{\text{max}}}$. During evaluation, the model is able to process studies containing up to $M_{\text{max}}$ scans without the need for interpolation. We investigate the effects of $M$ and $M_{\text{max}}$ in Figure 4(b), but we believe these hyperparameters should be tuned according to the distribution of the dataset.
@@ Appendix B Implementation
In this section, we summarize the model configuration described in Section 3.2 and provide details of the pre-training setup referenced in both Section 3.2 and Section 4.1. For linear probing evaluation on head CT, we refer readers to FM-HeadCT [zhu20253d].
@@ B.1 Chest CT
Visual Encoder ViT-B [dosovitskiy2020image]
@@
Text Encoder CXR-BERT [boecking2022making]
context length 512 [hamamci2024developing]
@@
optimizer AdamW [loshchilov2017decoupled]
@@
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.98 [radford2021learning]
learning rate schedule cosine decay [loshchilov2016sgdr]
warmup (in steps) 47 [shui2025large]
@@
CT-RATE [hamamci2024developing] Rad-ChestCT[draelos2021machine]
@@
The model architecture and pre-training configuration for chest CT are summarized in Table 4 and Table 4, respectively. These configurations closely follows prior works such as CT-CLIP [hamamci2024developing], OpenCLIP [cherti2023reproducible], and BiomedCLIP [zhang2023biomedclip]. In the pre-training configuration, we report the base learning rate corresponding to a batch size of 64. During training, we follow the linear learning rate scaling rule [li2023scaling]: $lr=base\_lr\times\frac{batch\_size}{64}$.
The diseases used for zero-shot evaluation are listed in Table 5. This setup is consistent with fVLM [shui2025large] and all methods reported in Table 1.
@@ B.2 Brain MRI/Head CT
Visual Encoder ViT-B [dosovitskiy2020image]
@@
Text Encoder PubMedBERT [pubmedbert, zhang2023biomedclip]
context length 256 [zhang2023biomedclip]
@@
optimizer AdamW [loshchilov2017decoupled]
@@
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.95 [li2023scaling]
learning rate schedule cosine decay [loshchilov2016sgdr]
@@
The model architecture and pre-training configuration for brain MRI/head CT are summarized in Table 7 and Table 7, respectively. These configurations also follows prior works [zhang2023biomedclip, cherti2023reproducible, li2023scaling]. In the pre-training configuration, we report the base learning rate corresponding to a batch size of 128. During training, we follow the linear learning rate scaling rule [li2023scaling]: $lr=base\_lr\times\frac{batch\_size}{128}$.
Appendix C Additional Ablation Studies on Chest CT
Table 8: Ablation studies of training differences between CT-CLIP [hamamci2024developing] and HLIP on the CT-RATE [hamamci2024developing] test split for internal validation and the Rad-ChestCT [draelos2021machine] for external validation.
AUC ACC F1 AUC ACC F1
CT-CLIP [hamamci2024developing] 73.3 66.9 70.8 63.3 59.9 64.7
ViT (our impl; bs=64) ${\texttt{+}2.0}$ ${\texttt{+}2.8}$ ${\texttt{+}2.4}$${\texttt{-}0.8}$${\texttt{-}0.8}$${\texttt{-}0.6}$
+ IN.MAE ${\texttt{+}0.8}$ ${\texttt{+}0.1}$ ${\texttt{+}0.2}$ ${\texttt{+}7.3}$ ${\texttt{+}6.3}$ ${\texttt{+}5.4}$
+ lock text ${\texttt{+}0.4}$ ${\texttt{+}0.5}$ ${\texttt{+}0.3}$ ${\texttt{+}1.2}$ ${\texttt{+}0.3}$ ${\texttt{+}0.2}$
+ hierarchical attn ${\texttt{+}0.3}$ ${\texttt{+}1.0}$ ${\texttt{+}0.8}$ ${\texttt{+}0.7}$ ${\texttt{+}1.1}$ ${\texttt{+}1.0}$
+batch size 512 +0.9 +0.1 +0.2 +0.6 +0.6 +1.4
= HLIP 77.7 71.4 74.7 72.3 68.4 72.1
[…]
$\rightarrow$[-1000,1000] 77.6 71.1 74.4 71.4 66.9 70.8
In this section, we address the question: What factors could influence language-image pre-training in radiology? under the constraint of a fixed data scale. Specifically, we evaluate the impact of several components added to the CT-CLIP baseline [hamamci2024developing] on chest CT benchmarks [draelos2021machine, hamamci2024developing]. As shown in Table 8:
  • ViT. Our implementation of the original ViT [dosovitskiy2020image] overfits well on the CT-RATE [hamamci2024developing] training set. Compared with CT-CLIP [hamamci2024developing], our implementation achieves a +2% AUC gain in internal evaluation and a -0.8% AUC drop in external evaluation. This can be attributed to the effective implementation adopted from OpenCLIP [cherti2023reproducible].
  • +MAE. MAE [he2022masked] pre-trained weight improves performance in external evaluation by a substantial margin of +7.3% AUC. This result highlights the universal features learned by MAE, as well as the advantage of using the original ViT backbone.
  • +Lock Text Encoder. During implementation, we observed severe overfitting even in the internal evaluation. Although the visual encoder is pre-trained with MAE, there exists a significant domain gap. In contrast, the text encoder is CXR-BERT [boecking2022making], whose training domain, chest X-ray diagnosis, substantially overlaps with chest CT. We hypothesize that the overfitting is primarily caused by the text encoder and thus freeze it during training. This strategy alleviates the overfitting issue and yields a +1.2% improvement in AUC.
  • HLIP (bs=512). Our HLIP further improves performance in both internal and external evaluations, achieving over 1% improvement across all three metrics, benefiting from both effective modeling and computational efficiency.
It is important to note that the performance improvement brought by HLIP being smaller than that of the MAE pre-trained weights does not imply that HLIP’s advantage lies solely in using MAE initialization. fVLM [shui2025large] also incorporates MAE pre-trained weights but still lags behind HLIP in both internal and external evaluations, as shown in Section 4.1. Therefore, it is the effectiveness of the complete HLIP framework that accounts for the observed improvement.
We additionally evaluate other factors that may be of interest to researchers during implementation. Among these, the most informative is saving the data in the uint8 format. Since language-image pre-training in 3D medical imaging requires loading volumetric data, data loading I/O can become a training-time bottleneck. As storing data in uint8 format does not affect performance, it offers a more efficient solution for both data loading and storage. Other factors, such as supervision summarized by Qwen [bai2023qwen], using resize operation instead of standardizing the spacing when constructing the batch, or different HU value truncation, only lead to subtle variations in internal evaluation. In external evaluation, we believe the performance variation is primarily caused by overfitting, which is expected to diminish as training scales up.
Therefore, identifying an effective strategy to scale up language-image pre-training in 3D medical imaging may be the most critical factor. Other design choices should be adopted insofar as they facilitate data scalability.
Appendix D Prospective Evaluation
@@
We present the detailed results of the prospective evaluation conducted within our health system for 52 brain MRI diagnoses and 83 head CT diagnoses in Figure 12 and Figure 13, respectively. We report linear probing AUC for HLIP (blue) and the original ViT model (red). HLIP demonstrate consistent improvement.
Appendix E Complexity
E.1 Analysis
@@
E.2 Experiments
We conduct experiments comparing ViT [dosovitskiy2020image], the 3D Swin Transformer implemented in MONAI [cardoso2022monai], and our HLIP visual encoder. Since Swin is not compatible with multi-scan inputs, the experiment is conducted on a single-channel 3D input of shape (B,1,224,224,224), where B=1 denotes the batch size. Both ViT and HLIP use a ViT-Base architecture with patch_size=(16,16,16). The Swin Transformer is configured with dim=128, patch_size=(4,4,4), window_size=(7,7,7), depth=(2,2,12,2), and head=(4,8,16,32). Following the original Swin Transformer [liu2021swin] while matching the model size.
@@ E.2 Experiments
w/o flash attn w/ flash attn [dao2022flashattention]
model size (M) throughput (img/s) $\uparrow$=better memory (G) $\downarrow$=better model size (M) throughput (img/s) $\uparrow$=better memory (G) $\downarrow$=better
ViT 88.2 9.0 6.6 88.2 44.4 1.5
Swin 87.9 9.9 8.0 87.9 9.9 9.0
HLIP (ours) 88.2 17.5 4.1 88.2 47.6 1.5
@@
All experiments are conducted on a single A40 GPU. We report the model size, throughput and training memory for each model. To measure throughput, we include 100 warm-up iterations and compute the average over 1000 forward passes. As shown in Table 9, when not using flash attention [dao2022flashattention], Swin is slightly faster than the original ViT but still lags behind our HLIP visual encoder. Swin also consumes more memory during training due to the large codebook of relative position embeddings constructed for 3D inputs. Moreover, Swin is not compatible with flash attention, which has been integrated into PyTorch 2.0 and supports a wide range of GPUs. We acknowledge that when using flash attention, the memory efficiency advantage of our hierarchical attention is reduced, since we do not alter the actual feature map size. However, HLIP remains faster than the original ViT.
Note that the results shown in Table 9 differ from those in Figure 1, which are measured on uncurated studies with input size (B,8,1,48,224,224) and without flash attention.
Appendix F Discussion
We propose HLIP, a scalable language-image pre-training framework for 3D medical imaging. In this section, we provide additional discussion of our approach.
Efficient Implementation. While we do not explicitly emphasize implementation efficiency in the paper, it is reflected in the training times reported in Section 3.2 and Section 4.1. To the best of our knowledge, training for 20 epochs on CT-RATE [hamamci2024developing] within 6 hours using 4 A40 GPUs represents the most efficient implementations in this domain. To support further development in language-image pre-training for 3D medical imaging, we will release our implementation soon.
Limitations. Here, we further discuss three limitations of our work:
  • The first limitation is computational resources. The most intensive compute setup used in this work is a node with 8 L40 GPUs, which is comparable to configurations used in other institutions. Although HLIP demonstrates substantial efficiency compared to the original ViT, we rely on both flash attention and gradient checkpointing to further optimize training. Nevertheless, the batch sizes achievable with our current resources are 256 for uncurated datasets and 512 for curated datasets which are still significantly smaller than the typical batch sizes used in the natural image and 2D medical imaging domains. Furthermore, our compute resources are not sufficient to train larger models such as ViT-Large.
  • We also observe that zero-shot transfer performance does not always correlate with the number of studies. For example, although keyword search yields more patients with meningioma or metastasis than glioma, the zero-shot performance for glioma is substantially higher than for the other two. While tumor size may partially explain this discrepancy, we believe it warrants further investigation in future work.

### v2

@@ 4.5 Visualizations
Figure 5: Qualitative results of zero-shot diagnosis on the Rad-ChestCT [24] and BraTS [47, 48] datasets. The first row shows the original clinical scans with pathologic regions outlined (dashed red). The second row shows activation HLIP maps [58]. HLIP identifies the pathologic regions across multiple groups of adjacent chest CT slides (left) and brain MRI scans (right).
@@
In Figure 5, we visualize the activation maps [58] of HLIP in the zero-shot setting on Rad-ChestCT [24] and BraTS [47, 48]. For chest CT (curated, single scan), HLIP attends to pathologic regions across different slices, e.g., slice Z and Z+8, by leveraging scan attention. For brain MRI (uncurated, multiple scans), HLIP attends to pathologic regions across different scan types, e.g., FLAIR and T1W+contrast, by leveraging study attention. This demonstrates the effectiveness HLIP in modeling the hierarchical structure of 3D medical imaging.
4.6 Clinical Translation of HLIP
@@
Figure 6: Prospective testing results on (a) 52 brain MRI diagnoses and (b) 83 head CT diagnoses. Results for each diagnosis are presented in Figures 11 and 12 in Appendix C.
We conduct a 1-year, health system-scale, comprehensive evaluation on a prospective set of neurological studies. The datasets include ˜23K brain MRI studies covering 52 diagnoses and ˜15K head CT studies covering 83 diagnoses. We report macro AUC in Figure 6(a) and 6(b). HLIP consistently outperforms ViT. We provide detailed results in Appendix C.
@@ 5 Conclusion
In this work, we pioneer language-image pre-training directly on uncurated 3D medical imaging studies containing multiple scans, together with a novel hierarchical attention mechanism that effectively extracts visual features. These two contributions form HLIP, a scalable and effective pre-training framework for 3D medical imaging. Despite its simplicity in both concept and practice, HLIP achieves state-of-the-art performance on multiple benchmarks across diverse modalities and anatomical regions, including brain MRI, head CT, and chest CT. These results demonstrate that directly pre-training on uncurated clinical datasets is a scalable and effective paradigm for language-image pre-training in 3D medical imaging. We hope our work facilitates large-scale language-image pre-training in other health systems and inspires future machine learning research on scalable approaches for real-world clinical data.
@@
  • C: Prospective Evaluation
  • D: Complexity
  • E: Discussion
@@ Appendix A Dataset
As mentioned in Section 3.2, this section provides additional details on the two datasets we collected: BrainMRI220K and HeadCT240K. For other datasets, such as CT-RATE [3], Rad-ChestCT [24], BraTS 2023 [59, 60, 47, 48, 49, 50], NYU-Mets [52], UCSF-Mets [51], RSNA [22], and CQ500 [21], we refer readers to their original publications. Additionally, we include a discussion of our preprocessing strategy.
@@ A.3 Discussion
To the best of our knowledge, this is the first work to handle 3D medical datasets at the scale of BrainMRI220K and HeadCT240K, which comprise millions of scans. Our preprocessing strategy, introduced in Section 3.2, follows standard practices where possible, but also diverges due to the scale and diversity of our dataset. Specifically, percentile clipping for brain MRI and HU value truncation for head CT follow standard practices [61, 20]. For head CT, we do not concatenate scans with different HU truncations along the channel dimension, as our method is capable of processing multiple scans simultaneously. However, our strategy regarding orientation and spacing may differ from established preprocessing pipelines [62, 4, 3, 6]. Here, we provide a discussion to justify the rationale behind our approach.
@@
Orientation. CT and MRI scans are acquired using different planes, including axial, coronal, and sagittal. Isotropic data are rare in routine radiology studies; the through-plane spacing is typically larger than the in-plane spacing. In the context of batch construction, where all samples in a batch must share the same shape, standardizing the orientation can lead to downsampling along the in-plane axes and upsampling along the through-plane axis. Upsampling in pixel space does not introduce new information, whereas downsampling results in information loss. From the perspective of data augmentation, the orientation can change when 3D random rotation is applied [62]. So, why do we standardize the orientation at the beginning then augment it during training? Instead, we consistently align the first dimension (depth) with the through-plane axis. Please note that the in-plane orientation is still standardized as RP, RI, or PI. We treat the diversity of acquisition planes (as shown in Figure 7(c) and Figure 9(c)) in our dataset as a form of natural data augmentation.
@@
Spacing. Standardizing spacing across the entire dataset is a common practice in medical image segmentation to ensure that convolution filters interpret anatomical structures consistently [62]. However, HyperSpace [63] demonstrates that, with appropriate spacing augmentation, convolutional networks can learn spacing variations. Therefore, architectures like Vision Transformers, which can learn scale-invariant features [11], should also be capable of learning spacing-invariant features with sufficient spacing augmentation. This is because spacing invariance is a direct extension of scale equivariance from 2D to 3D space. Given the diversity of spacing settings (as shown in Figure 7(b) and Figure 9(b)) in our dataset, we believe it naturally occupies sufficient spacing augmentation for the model to learn spacing-equivariant features. Therefore, when constructing a batch, we first resize each volume to (48,256,256), and then apply a center crop to obtain (48,224,224). Here, 48 represents the median number of slices, which does not substantially distort the original scan.
@@
$M$ scans per study. As the number of scans per study varies, we randomly sample $M=10$ scans per study at each training step to construct a batch. An equal number of positional embeddings is also sampled for these $M$ scans from a total of $M_{\text{max}}=40$ positional embeddings at each step. This means the model is exposed to $M_{\text{max}}$ different positions over the course of training, while observing $M$ positions at each step. This is analogous to a dropout strategy, with a scan dropout rate of $1-\frac{M}{M_{\text{max}}}$. During evaluation, the model is able to process studies containing up to $M_{\text{max}}$ scans without the need for interpolation.
@@ Appendix B Implementation
In this section, we summarize the model configuration described in Section 3.2 and provide details of the pre-training setup referenced in both Section 3.2 and Section 4.3. For linear probing evaluation on head CT, we refer readers to FM-HeadCT [20].
@@ B.1 Chest CT
Visual Encoder ViT-B [11]
@@
Text Encoder CXR-BERT [7]
context length 512 [3]
@@
optimizer AdamW [64]
@@
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.98 [2]
learning rate schedule cosine decay [65]
warmup (in steps) 47 [6]
@@
CT-RATE [3] Rad-ChestCT [24]
@@
The model architecture and pre-training configuration for chest CT are summarized in Table 4 and Table 4, respectively. These configurations closely follows prior works such as CT-CLIP [3], OpenCLIP [45], and BiomedCLIP [32]. In the pre-training configuration, we report the base learning rate corresponding to a batch size of 64. During training, we follow the linear learning rate scaling rule [19]: $lr=base\_lr\times\frac{batch\_size}{64}$.
The diseases used for zero-shot evaluation are listed in Table 5. This setup is consistent with fVLM [6] and all methods reported in Table 2.
@@ B.2 Brain MRI/Head CT
Visual Encoder ViT-B [11]
@@
Text Encoder PubMedBERT [44, 32]
context length 256 [32]
@@
optimizer AdamW [64]
@@
optimizer momentum $\beta_{1}$, $\beta_{2}$ = 0.9, 0.95 [19]
learning rate schedule cosine decay [65]
@@
The model architecture and pre-training configuration for brain MRI/head CT are summarized in Table 7 and Table 7, respectively. These configurations also follows prior works [32, 45, 19]. In the pre-training configuration, we report the base learning rate corresponding to a batch size of 128. During training, we follow the linear learning rate scaling rule [19]: $lr=base\_lr\times\frac{batch\_size}{128}$.
Appendix C Prospective Evaluation
@@
We present the detailed results of the prospective evaluation conducted within our health system for 52 brain MRI diagnoses and 83 head CT diagnoses in Figure 11 and Figure 12, respectively. We report linear-probe AUC for HLIP (blue) and the vanilla ViT (red). HLIP demonstrate consistent improvement.
Appendix D Complexity
D.1 Analysis
@@
D.2 Experiments
We conduct experiments comparing ViT [11], the 3D Swin Transformer implemented in MONAI [66], and our HLIP visual encoder. Since Swin is not compatible with multi-scan inputs, the experiment is conducted on a single-channel 3D input of shape (B,1,224,224,224), where B=1 denotes the batch size. Both ViT and HLIP use a ViT-Base architecture with patch_size=(16,16,16). The Swin Transformer is configured with dim=128, patch_size=(4,4,4), window_size=(7,7,7), depth=(2,2,12,2), and head=(4,8,16,32). Following the original Swin Transformer [12] while matching the model size.
@@ D.2 Experiments
model size
throughput
(img/s)
$\uparrow$=better
memory
$\downarrow$=better
ViT 88.2 9.0 6.6
Swin 87.9 9.9 8.0
HLIP (ours) 88.2 17.5 4.1
@@
All experiments are conducted on a single A40 GPU. We report the model size, throughput and training memory for each model. To measure throughput, we include 100 warm-up iterations and compute the average over 1000 forward passes. As shown in Table 8, Swin is slightly faster than the original ViT but lags behind our HLIP visual encoder. Swin also consumes more memory during training due to the large codebook of relative position embeddings constructed for 3D inputs. Moreover, Swin is not compatible with flash attention, which has been integrated into PyTorch 2.0 and supports a wide range of GPUs.
Appendix E Discussion
In this work, we present HLIP, a scalable language–image pre-training framework for 3D medical imaging. For health systems that have accumulated large volumes of data over the past decades, HLIP offers a new direction for learning transferable representations. In this section, we provide additional discussion, including both the limitations and future directions for readers interested in pursuing this line of research.